OpenStudy (anonymous):

Integrate sqrtx / sqrtx + 1

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ \sqrt{x} }{ \sqrt{x} + 1 }dx\]

OpenStudy (anonymous):

I'ved tried substitution, but it gets more complex

hartnn (hartnn):

multiply and divide by root x-1

OpenStudy (anonymous):

sorry?

OpenStudy (anonymous):

rationalize the denominator

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ \sqrt{x} }{ \sqrt{x} + 1 } * \frac{ \sqrt{x} + 1 }{ \sqrt{x} + 1 } dx\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

hmm,ok

OpenStudy (anonymous):

teacher never showed us this technique

OpenStudy (anonymous):

then just do substitution afterwards?

OpenStudy (anonymous):

it gets more complex

hartnn (hartnn):

oh, wait...that won't work....

hartnn (hartnn):

put u = root x +1

OpenStudy (anonymous):

oo maybe that'll work

hartnn (hartnn):

then root x = u-1

hartnn (hartnn):

dx= 2(u-1) du

hartnn (hartnn):

and then it becomes easy to solve ..... :)

hartnn (hartnn):

ask if any doubts or if u didn't get it...

OpenStudy (anonymous):

ok, ill post my answer in a sec

OpenStudy (anonymous):

\[\int\limits_{}^{}2u + \int\limits_{}^{}\frac{ 2 }{ u }-\int\limits_{}^{}4 \]

hartnn (hartnn):

yes, going good....proceed.

OpenStudy (anonymous):

is the antiderivative of 4, 4x or 4u

OpenStudy (anonymous):

4u cause im in terms of u right now

hartnn (hartnn):

yes, 4u

OpenStudy (anonymous):

\[u^2+u-4u\] \[(\sqrt{x}+1)^2+(\sqrt{x}+1)-4(\sqrt{x}+1)\]

hartnn (hartnn):

for 2/u , its 2log u

OpenStudy (anonymous):

oh yea..

OpenStudy (anonymous):

\[(\sqrt{x}+1)^2+2\ln(\sqrt{x}+1)-4(\sqrt{x}+1)\]

OpenStudy (anonymous):

then expand

hartnn (hartnn):

correct.

OpenStudy (anonymous):

ok thanks hartnn!

hartnn (hartnn):

welcome ^_^